OpenStudy (anonymous):

Choose the equation below that represents the line passing through the point (-2, -3) with a slope of -6. a. y = -6x - 15 b. y = -6x - 20 c. y = -6x + 15 d. y = -6x + 20

OpenStudy (anonymous):

The equation we'll be using here is: \[y-y_1=m(x-x_1)\]Where m is the slope of the line, and x_1,y_1 are the coordinates of any point on the line. We have the slope and a point on the line, so you can plug in and solve.

OpenStudy (anonymous):

Can you plug choice a into that equation so I can understand how to use it please? I have several of these to do.

OpenStudy (anonymous):

vinnv226 I don't understand which numbers to plug in

OpenStudy (anonymous):

x1 = -2 y1= -3 m = -6 \[y + 3 = -6(x+2)\] solve for y :)

OpenStudy (anonymous):

m always represents the slope of the line. The other variables, x_1 and y_1, are the x and y coordinates of any point on the line, as @hobbs978 showed above

OpenStudy (anonymous):

y = -6(x+2-3) y = -6x -1 but that isn't one of the choices. I'm so confused!

OpenStudy (anonymous):

You have to distribute the -6 through the parentheses before you can subtract the 3 over. You can't just subtract 3 from the inside of the parentheses.

OpenStudy (anonymous):

so the correct answer is c. y = -6x + 15

OpenStudy (anonymous):

so the correct answer is c. y = -6x + 15?

OpenStudy (anonymous):

No, thats not correct. Once you distribute the -6 you'll get this: y+3=-6x -12

OpenStudy (anonymous):

shoot!

OpenStudy (anonymous):